# A multi-band study of pulsar glitches with Fermi-LAT and Parkes

### v1

@@ 4 RESULTS
Table 3 reports the parameters of all the detected glitches. The first three columns are the pulsar name, the reference number of each glitch, and the glitch epoch. Glitch epoch is obtained by extrapolation where the phase continuity was lost in the timing solution. It is necessary to use Eq. (1) to fit the timing solutions of pre-glitch and post-glitch respectively. Glitch epoch is the middle of interval between the last pre-glitch observation time and the first post-glitch observation time. The uncertainty is half of the interval between the two observations. The other parameters are obtained through phase-coherent timing fits, that is, fitting Eq. (2) to ToAs, while the associated uncertainty is calculated using the error transfer equation. The fourth and fifth columns show the relative change in the spin frequency and the first derivative of frequency, respectively. In the sixth column, Y represents the new detected glitch in this work and P is the previously published glitch event. If significant exponential recovery process is observed, we will add exponential decay term to the TEMPO2 fitting, the results of these parameters are shown in the seventh and eighth columns. The last two columns are the RMS and the data span. The results of the timing analysis are shown in Figs. (1), (2), (3), (5) and (6), where value of each $\nu$ and $\dot{\nu}$ was obtained through independent fitting of ToAs of 150–600 d. In this work, we use the least squares method to fit the exponential decay and linear recovery of the spin-down rate after the glitch, respectively. Specifically, the red curves in these figures represent the fitting results of exponential recovery, while the blue and magenta dotted curves represent the fitting results of two different linear recoveries.
Next, we will specifically describe the glitch events. The fits within the framework of the vortex creep model (see Sec. 3) are done for 9 large glitches in our sample, using Eqs. (7), (10) and (14). The results are reported along with Figs. (1), (2), (3), (5) and in Table 4. One sees that the recovery fit from a different and physically-motivated functional form describes the observations equally well. Various moments of inertia from the vortex-creep-model fit are collected in Table 5. The results suggest that glitches involve only the crustal superfluid and possibly a portion of the outer core, with the adopted creep model providing sufficient capacity for the observed glitch amplitudes. In the next section of Sec. 5, we also show our results for the oscillations in the spin-down rates of PSRs J1509--5850 and J1718--3825, and reassess them within the vortex bending model. Moreover, beyond the common magnetic dipole-radiation scenario, we discuss the effect of the internal torque from bending vortex and inter-glitch creep contributions on the braking index.
@@ 4.1 PSR J1420--6048
Figure 1: Glitches in PSR J1420--6048: The vertical gray dashed line and the number of parentheses at the top represent the glitch epoch and the reference number of glitch occurrences, respectively. Panel (a) shows the relative change ($\Delta\nu$) of spin frequency, that is, the change in pulsar rotation frequency relative to the frequency model before the glitch. Panel (b) is an enlarged view of panel (a), showing the evolution details of $\Delta\nu$ after the glitch, which is obtained by subtracting its average value from $\Delta\nu$ of post-glitch. Panel (c) shows the evolution of the spin-down rate ($\dot{\nu}$) over time before and after the glitch. The red, blue, and magenta dashed lines represent the one exponential and two different linear recovery fitting results, respectively. In Panel (d), the green curves represent the fit using the vortex creep model (see Sec. 3), with the corresponding fit parameters provided in Table 4.
PSR J1420--6048 is an energetic Vela-like pulsar, first discovered during the Parkes multilateral pulsar survey by D’Amico et al. (2001). In addition to being observed in the radio band, this young pulsar has been detected in both the X-ray and gamma-ray bands (Weltevrede et al., 2010). X-ray pulsations (AX J1420.1--6049) were observed by Roberts et al. (2001) using ASCA, while gamma-ray emissions (GeV J1417--6100/3EG J1420--6038) were detected by the EGRET telescope onboard the Compton Gamma Ray Observatory (Nishida et al., 2003).
@@
To date, eight glitches have been observed in PSR J1420--6048 (Yu et al., 2013; Liu et al., 2021; Lower et al., 2021; Lin et al., 2021), with fractional glitch sizes on the order of $\Delta\nu/\nu\sim 10^{-6}$. These glitches generally exhibit exponential recoveries, with glitch intervals of 2–3 yr. Lin et al. (2021) found that in the gamma-ray band, the pulse profile of PSR J1420--6048 may change significantly over different time spans.
@@
We have combined Fermi-LAT data with Parkes radio timing data to update the glitch epochs and parameters for glitches 4, 5, 6, 7, and 8, as shown in Table 3. Fig. 1 shows the 14 yr evolution of $\nu$ and $\dot{\nu}$ for PSR J1420--6048. Previous studies indicated no exponential decay component in the post-glitch recovery of glitch 8 (Lin et al., 2021). However, our analysis of the fitted timing residuals reveals a clear exponential recovery component, as shown in Fig. 1. We modeled the post-glitch recovery for glitch 8, finding an exponential timescale of 100(6) d and $Q=0.0131(5)$. Additionally, we updated the fractional glitch size and its first derivative to $\Delta\nu/\nu\sim 1480(1)\times 10^{-9}$ and $\Delta\dot{\nu}/\dot{\nu}\sim 6.0(1)\times 10^{-3}$, respectively.
@@
From panel (c) of Fig. 1, we observe that glitches 4, 6, and 8 each exhibit two distinct linear recovery phases following an initial exponential recovery. This is particularly evident for glitch 4. To verify this, we performed exponential and linear fits to the evolution of $\dot{\nu}$. In panel (c), the red curve represents the exponential fit, while the blue and magenta dashed lines correspond to the first and second linear recovery phases, respectively. Just as shown in panel (c), the linear processes of glitch 4 and glitches 6, 8 are clearly different. For glitch 4, the slope of the first linear recovery phase ($\ddot{\nu}\sim 1.778\times 10^{-21}$ s-3) is nearly double that of the second phase ($\ddot{\nu}\sim 8.469\times 10^{-22}$ s-3). In contrast, glitches 6 and 8 show the opposite trend: the slopes of the second linear phase are larger than the first. Specifically, glitch 6 has $\ddot{\nu}$ values of $5.454\times 10^{-22}$ s-3 and $8.364\times 10^{-22}$ s-3, while glitch 8 exhibits values of $8.388\times 10^{-22}$ s-3 and $1.237\times 10^{-21}$ s-3 for the two linear phases, respectively. For glitches 5 and 7, a single linear recovery phase adequately describes the post-glitch evolution, with slopes of $1.185\times 10^{-21}$ s-3 and $6.327\times 10^{-22}$ s-3, respectively. This behaviour of two consecutive linear recovery processes in the post-glitch has not been previously reported for PSR J1420--6048, likely due to the limited timing data available from individual Parkes and Fermi-LAT observations.  [^6] Nevertheless, it should be mentioned here that the “shoulder” or “elongated-S” shape seen after the glitches in PSR J1420--6048 may also be due to an unaccounted part of the post-glitch recovery times which could be originated from unmodelled timing noise as discussed in (Antonopoulou et al., 2022). More discussions on recovery behaviour can be found below in Sec. 5.
@@
  • G4: We fit the fourth glitch (G4) of PSR J1420--6048 with Eqs. (7) and (10). As seen in Table 4, the vortex creep model prediction for the time to the next glitch from G4 is $t_{\rm th}=698(32)$ d which is in qualitative agreement with the observed inter-glitch time $t_{\rm obs}=747(28)$ d within uncertainties. The exponential decay time-scale obtained from the vortex creep model application $\tau_{\rm creep,exp}=41(11)$ days is slightly different than what found from fit with Eq. (2): $\tau_{\rm d}=53(7)$ d (see Table 3). The reason is that Eq. (10) has also an exponential dependence and dominates over the initial days of the data. The recoupling time-scale for the crustal superfluid associated with G4 is $\tau_{\rm s}=116(10)$ d. The total fractional crustal superfluid moment of inertia involved in G4 is $I_{\rm cs}/I=2.56(6)\%$.
@@
  • G5: We fit the fifth glitch (G5) of PSR J1420--6048 with Eqs. (10) and (14). Due to the initial exponential dependencies of the two equations, there is no need to include a separate exponential function to fit the data of G5. As seen in Table 4, the vortex creep model prediction for the time to the next glitch from G5 is $t_{\rm th}=964(19)$ d, while the observed inter-glitch time $t_{\rm obs}=856(10)$ d. The origin of the discrepancy may lie in the fact that crustquake, which played the role of starting the vortex unpinning avalanche, probably left behind a tiny persistent shift in the spin-down rate by leading to a horizontal displacement of the broken platelet. Such effects on the vortex creep model estimates are also visible for some of the Crab (Gügercinoǧlu & Alpar, 2019) and the Vela (Akbal et al., 2017) glitches. The recoupling time-scale for the crustal superfluid associated with G5 is $\tau_{\rm s}=109(6)$ d. The total fractional crustal superfluid moment of inertia involved in G5 is $I_{\rm cs}/I=2.34(34)\%$. Both $\tau_{\rm s}$ and $I_{\rm cs}/I$ are quite similar for G4 and G5, suggesting that these two glitches have originated from the same crustal superfluid region.
@@
  • G6: We fit the sixth glitch (G6) of PSR J1420--6048 with Eqs. (7), (10) and (14). As seen in Table 4, the exponential decay time-scale obtained from the vortex creep model application $\tau_{\rm creep,exp}=84(11)$ d is slightly different than $\tau_{\rm d}=44(8)$ d found from fit with Eq. (2) (see Table 3) again due to the initial exponential dependencies of Eqs. (10) and (14). The vortex creep model prediction for the time to the next glitch from G6 is $t_{\rm th}=1124(32)$ d, while the observed inter-glitch time is $t_{\rm obs}=960(13)$ d. However, as can be clearly seen from Fig. 1, the seventh glitch (G7) arrived before the recovery of G6 has been fully completed. The recoupling time-scale for the crustal superfluid associated with G6 is $\tau_{\rm s}=98(11)$ d, similar to the value of the previous two events. The total fractional crustal superfluid moment of inertia involved in G6 is $I_{\rm cs}/I=1.54(2)\%$.
@@
  • G7: We fit the seventh glitch (G7) of PSR J1420--6048 with Eqs. (7), and (14). As seen in Table 4, the exponential decay time-scale obtained from the vortex creep model application $\tau_{\rm creep,exp}=109(32)$ d is consistent with $\tau_{\rm d}=102(7)$ d found from fit with Eq. (2) (see Table 3). As can be clearly seen from Fig. 1, the post-glitch spin-down rate of G7 over-recoveries, i.e. extends over the original pre-glitch level. One possible explanation of this peculiar and rare type of relaxation in terms of the vortex creep model is as follows. According to the vortex creep model, there are two nonlinear creep regions responsible for the post-glitch rotation and spin-down rate evolution, namely vortex traps (initiating a glitch through collective vortex unpinning) with moment of inertia $I_{\rm A}$ and vortex capacitor regions (inside of which vortices cannot creep but other freed lines may pass through of it and take part in angular momentum conservation) with moment of inertia $I_{\rm B}$. If some of these capacitor regions become activated such that vortices inside that region start to creep after their interaction with the unpinned ones, then with the excess angular momentum transfer of such activated vortices the post-glitch spin-down rate will extend over the pre-glitch level. The vortex creep model prediction for the time to the next glitch from G7 is $t_{\rm th}=938(425)$ d and agrees with the observed inter-glitch time $t_{\rm obs}=1339(14)$ d within errors. The recoupling time-scale for the crustal superfluid associated with G6 is $\tau_{\rm nl}=779(554)$ d and due to its being longer has not affected the initial exponential decay part of the glitch. The total fractional crustal superfluid moment of inertia involved in G7 is $I_{\rm cs}/I=2.05(1.98)\%$.
@@
  • G8: We fit the eighth glitch (G8) of PSR J1420--6048 with Eqs (7), (10) and (14). As seen in Table 4, the exponential decay time-scale obtained from the vortex creep model application $\tau_{\rm creep,exp}=86(39)$ d is slightly different than $\tau_{\rm d}=100(6)$ d found from fit with Eq. (2) (see Table 3) again due to the initial exponential dependencies of Eqs. (10) and (14). The recoupling time-scale for the crustal superfluid associated with G8 is $\tau_{\rm s}=26(11)$ d and smallest among the glitches of PSR J1048--6048 analysed in this study. This value is consistent with its lower superfluid reservoir involvement in the glitch. The total fractional crustal superfluid moment of inertia involved in G8 is $I_{\rm cs}/I=1.23(2)\%$.

### v2

@@ 4 RESULTS
Table 3 reports the parameters of all the detected glitches. The first three columns are the pulsar name, the reference number of each glitch, and the glitch epoch. The middle of interval between the last pre-glitch observation time and the first post-glitch observation time is determined as the glitch epoch. The uncertainty is half of the interval between the two observations. The other parameters are obtained through fitting Eq. (2) to ToAs, while the associated uncertainty is calculated using the error propagation. Specifically, the fourth and fifth columns show the relative change in the spin frequency ($\Delta\nu/\nu$) and the first derivative of frequency ($\Delta\dot{\nu}/\dot{\nu}$), respectively. In the sixth column, Y represents the newly detected glitch in this work and P is the previously published glitch event. If significant exponential recovery process is observed, we added exponential decay term to the TEMPO2 fitting, the results of these parameters ($Q$, $\tau_{\rm d}$) are shown in the seventh and eighth columns. The last two columns are the RMS and the data span.
Besides, the evolutions of $\nu$ and $\dot{\nu}$ are shown in Figs. (1-3) and (5-6), where values of each $\nu$ and $\dot{\nu}$ were obtained through independent fitting Eq. (1) to ToAs data spanning 150–600 d. Moreover, we applied the least squares method to perform piecewise fitting for $\dot{\nu}$ data, using the reduced chi-square value ($\chi^{2}_{\rm r}$) to evaluate the quality of the fit. Exponential fitting was applied to the $\dot{\nu}$ data within a few hundred days after the glitch, while linear fitting was applied to the $\dot{\nu}$ data following the exponential decay.
In the panels (d) of Figs. (1-3) and (5) , we also report the fits within the framework of the vortex creep model (see Sec. 3) (shown in green curves), done for 9 large glitches in our sample, using Eqs. (7), (10) and (14). The corresponding fit parameters are provided in Table 4. Based on the $\chi^{2}_{\rm r}$ value of the fits, one sees that the recovery fit from a different and physically-motivated functional form describes the observations equally well. Besides, various moments of inertia from the vortex-creep-model fit are collected in Table 5. The results suggest that glitches involve only the crustal superfluid and possibly a portion of the outer core, with the adopted creep model providing sufficient capacity for the observed glitch amplitudes.
In the next section of Sec. 5, we also show our results for the oscillations in the spin-down rates of PSRs J1718--3825 and J1509--5850, and reassess them within the vortex bending model (see purple curves in Figs. 5-6). Moreover, beyond the common magnetic dipole-radiation scenario, we discuss the effect of the internal torque from bending vortex and inter-glitch creep contributions on the braking index ($n=\nu\ddot{\nu}/\dot{\nu}^{2}$).
@@ 4.1 PSR J1420--6048
Figure 1: Glitches in PSR J1420--6048: The vertical gray dashed line and the numbers in between parentheses at the top represent the glitch epoch and the reference number of glitch occurrences, respectively. Panel (a) shows the relative change ($\Delta\nu$) of spin frequency, that is, the change in pulsar rotation frequency relative to the frequency model before the glitch. Panel (b) is an enlarged view of panel (a), showing the evolution details of $\Delta\nu$ after the glitch, which is obtained by subtracting its average value from $\Delta\nu$ of post-glitch. Panel (c) shows the evolution of the spin-down rate ($\dot{\nu}$) over time before and after the glitch. The red, blue, and magenta dashed lines represent the one exponential and two different linear recovery fitting results, respectively. In Panel (d), the green curves represent the fit using the vortex creep model (see Sec. 3), with the corresponding fit parameters provided in Table 4.
PSR J1420--6048 is an energetic Vela-like pulsar, first discovered during the Parkes multibeam pulsar survey by D’Amico et al. (2001). In addition to being observed in the radio band, this young pulsar has been detected in both the X-ray and gamma-ray bands (Weltevrede et al., 2010). X-ray pulsations (AX J1420.1--6049) were observed by Roberts et al. (2001) using ASCA, while gamma-ray emissions (GeV J1417--6100/3EG J1420--6038) were detected by the EGRET telescope onboard the Compton Gamma Ray Observatory (Nishida et al., 2003).
@@
To date, eight glitches have been observed in PSR J1420--6048 (Yu et al., 2013; Liu et al., 2021; Lower et al., 2021; Lin et al., 2021), with fractional glitch sizes on the order of $\Delta\nu/\nu\sim 10^{-6}$. These glitches generally exhibit exponential recoveries, with glitch intervals of 2–3 yr. As mentioned above, Lin et al. (2021) found that in the gamma-ray band, the pulse profile of PSR J1420--6048 may change significantly before and after glitches 4 and 8.
@@
We have combined Fermi-LAT data with Parkes radio timing data to update the glitch epochs and parameters for glitches 4, 5, 6, 7, and 8, as shown in Table 3. Fig. 1 shows the 14 yr evolution of $\nu$ and $\dot{\nu}$ for PSR J1420--6048. Previous studies indicated no exponential decay component in the post-glitch recovery of glitch 8 (Lin et al., 2021). However, our analysis reveals a clear exponential recovery component, as shown in Fig. 1. We modeled the post-glitch recovery for glitch 8, finding an exponential timescale of 100(6) d and $Q=0.0131(5)$. Additionally, we updated the fractional glitch size and its first derivative to $\Delta\nu/\nu\sim 1480(1)\times 10^{-9}$ and $\Delta\dot{\nu}/\dot{\nu}\sim 6.0(1)\times 10^{-3}$, respectively.
@@
From panel (c) of Fig. 1, we observe that glitches 4, 6, and 8 each exhibit two distinct linear recovery phases following an initial exponential recovery. This is particularly evident for glitch 4. To verify this, we performed exponential and linear fits to the evolution of $\dot{\nu}$ (see panel (c)), and calculated the $\chi^{2}_{\rm r}$ values of these fits, which range from 0.57 to 2.25, with corresponding to degrees of freedom (d.o.f.) ranging from 2 to 11, respectively. Just as shown in panel (c), the linear processes of glitch 4 and glitches 6, 8 are clearly different. For glitch 4, the slope of the first linear recovery phase ($\ddot{\nu}\sim 1.778(3)\times 10^{-21}$ s-3) is nearly double that of the second phase ($\ddot{\nu}\sim 8.58(6)\times 10^{-22}$ s-3). In contrast, glitches 6 and 8 show the opposite trend: the slopes of the second linear phase are larger than the first. Specifically, glitch 6 has $\ddot{\nu}$ values of $5.45(2)\times 10^{-22}$ s-3 and $8.49(5)\times 10^{-22}$ s-3, while glitch 8 exhibits values of $8.39(2)\times 10^{-22}$ s-3 and $1.237(6)\times 10^{-21}$ s-3 for the two linear phases, respectively. For glitches 5 and 7, a single linear recovery phase adequately describes the post-glitch evolution, with slopes of $1.185(2)\times 10^{-21}$ s-3 and $6.327(9)\times 10^{-22}$ s-3, respectively. This behaviour of two consecutive linear recovery processes in the post-glitch has not been previously reported for PSR J1420--6048, likely due to the limited timing data available from individual Parkes and Fermi-LAT observations.  [^8] Nevertheless, it should be mentioned here that the “shoulder” or “elongated-S” shape seen after the glitches in PSR J1420--6048 may also be due to an unaccounted part of the post-glitch recovery times which could be originated from unmodelled timing noise as discussed in (Antonopoulou et al., 2022). More discussions on recovery behaviour can be found below in Sec. 5.
@@
  • G4: We fit the fourth glitch (G4) of PSR J1420--6048 with Eqs. (7) and (10), with a $\chi^{2}_{\rm r}$ value of 1.17 (12 d.o.f.). As seen in Table 4, the vortex creep model prediction for the time to the next glitch from G4 is $t_{\rm th}=698(32)$ d which is in qualitative agreement with the observed inter-glitch time $t_{\rm obs}=747(28)$ d within uncertainties. The exponential decay time-scale obtained from the vortex creep model application $\tau_{\rm creep,exp}=41(11)$ days is slightly different than what found from fit with Eq. (2): $\tau_{\rm d}=53(7)$ d (see Table 3). The reason is that Eq. (10) has also an exponential dependence and dominates over the initial days of the data. The recoupling time-scale for the crustal superfluid associated with G4 is $\tau_{\rm s}=116(10)$ d. The total fractional crustal superfluid moment of inertia involved in G4 is $I_{\rm cs}/I=2.56(6)\%$.
@@
  • G5: We fit the fifth glitch (G5) of PSR J1420--6048 with Eqs. (10) and (14), with the $\chi^{2}_{\rm r}$ value is 1.01 (13 d.o.f.). Due to the initial exponential dependencies of the two equations, there is no need to include a separate exponential function to fit the data of G5. As seen in Table 4, the vortex creep model prediction for the time to the next glitch from G5 is $t_{\rm th}=964(19)$ d, while the observed inter-glitch time $t_{\rm obs}=856(10)$ d. The origin of the discrepancy may lie in the fact that crustquake, which played the role of starting the vortex unpinning avalanche, probably left behind a tiny persistent shift in the spin-down rate by leading to a horizontal displacement of the broken platelet. Such effects on the vortex creep model estimates are also visible for some of the Crab (Gügercinoǧlu & Alpar, 2019) and the Vela (Akbal et al., 2017) glitches. The recoupling time-scale for the crustal superfluid associated with G5 is $\tau_{\rm s}=109(6)$ d. The total fractional crustal superfluid moment of inertia involved in G5 is $I_{\rm cs}/I=2.34(34)\%$. Both $\tau_{\rm s}$ and $I_{\rm cs}/I$ are quite similar for G4 and G5, suggesting that these two glitches have originated from the same crustal superfluid region.
@@
  • G6: We fit the sixth glitch (G6) of PSR J1420--6048 with Eqs. (7), (10) and (14), with the $\chi^{2}_{\rm r}$ value is 0.60 (12 d.o.f.). As seen in Table 4, the exponential decay time-scale obtained from the vortex creep model application $\tau_{\rm creep,exp}=84(11)$ d is slightly different than $\tau_{\rm d}=44(8)$ d found from fit with Eq. (2) (see Table 3) again due to the initial exponential dependencies of Eqs. (10) and (14). The vortex creep model prediction for the time to the next glitch from G6 is $t_{\rm th}=1124(32)$ d, while the observed inter-glitch time is $t_{\rm obs}=960(13)$ d. However, as can be clearly seen from Fig. 1, the seventh glitch (G7) arrived before the recovery of G6 has been fully completed. The recoupling time-scale for the crustal superfluid associated with G6 is $\tau_{\rm s}=98(11)$ d, similar to the value of the previous two events. The total fractional crustal superfluid moment of inertia involved in G6 is $I_{\rm cs}/I=1.54(2)\%$.
@@
  • G7: We fit the seventh glitch (G7) of PSR J1420--6048 with Eqs. (7), and (14), with very low $\chi^{2}_{\rm r}$ value (0.43, for 18 d.o.f.). As seen in Table 4, the exponential decay time-scale obtained from the vortex creep model application $\tau_{\rm creep,exp}=109(32)$ d is consistent with $\tau_{\rm d}=102(7)$ d found from fit with Eq. (2) (see Table 3). As can be clearly seen from Fig. 1, the post-glitch spin-down rate of G7 over-recoveries, i.e. extends over the original pre-glitch level. One possible explanation of this peculiar and rare type of relaxation in terms of the vortex creep model is as follows. According to the vortex creep model, there are two nonlinear creep regions responsible for the post-glitch rotation and spin-down rate evolution, namely vortex traps (initiating a glitch through collective vortex unpinning) with moment of inertia $I_{\rm A}$ and vortex capacitor regions (inside of which vortices cannot creep but other freed lines may pass through of it and take part in angular momentum conservation) with moment of inertia $I_{\rm B}$. If some of these capacitor regions become activated such that vortices inside that region start to creep after their interaction with the unpinned ones, then with the excess angular momentum transfer of such activated vortices the post-glitch spin-down rate will extend over the pre-glitch level. The vortex creep model prediction for the time to the next glitch from G7 is $t_{\rm th}=938(425)$ d and agrees with the observed inter-glitch time $t_{\rm obs}=1339(14)$ d within errors. The recoupling time-scale for the crustal superfluid associated with G6 is $\tau_{\rm nl}=779(554)$ d and due to its being longer has not affected the initial exponential decay part of the glitch. The total fractional crustal superfluid moment of inertia involved in G7 is $I_{\rm cs}/I=2.05(1.98)\%$.
@@
  • G8: We fit the eighth glitch (G8) of PSR J1420--6048 with Eqs (7), (10) and (14), and the corresponding $\chi^{2}_{\rm r}$ value of 1.87 (8 d.o.f.). As seen in Table 4, the exponential decay time-scale obtained from the vortex creep model application $\tau_{\rm creep,exp}=86(39)$ d is slightly different than $\tau_{\rm d}=100(6)$ d found from fit with Eq. (2) (see Table 3) again due to the initial exponential dependencies of Eqs. (10) and (14). The recoupling time-scale for the crustal superfluid associated with G8 is $\tau_{\rm s}=26(11)$ d and smallest among the glitches of PSR J1048--6048 analysed in this study. This value is consistent with its lower superfluid reservoir involvement in the glitch. The total fractional crustal superfluid moment of inertia involved in G8 is $I_{\rm cs}/I=1.23(2)\%$.